OpenStudy (anonymous):

How many and of what type are the solutions to x^2 + 10x + 25 = 0?

OpenStudy (anonymous):

X^2+10X+25=0 Because it is a quadratic eq. Hence it has two solutions. x^2+5x+5x+25=0 x(x+5)+5(x+5)=0 (x+5)(x+5)=0 x+5=0, gives x=-5 solution is x=-5,-5

OpenStudy (anonymous):

so that would be c

OpenStudy (anonymous):

@surjithayer